OpenStudy (anonymous):

what values for theta(0

OpenStudy (thomas5267):

Factor out \(\cos(\theta)\). Then solve the two brackets.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

There are 4 solutions.

OpenStudy (thomas5267):

So let there be four solutions.

OpenStudy (anonymous):

How do I do this?

OpenStudy (thomas5267):

\[ 2\cos(\theta)\sin(\theta)+\sqrt{3}\cos(\theta)=0\\ \cos(\theta)(2\sin(\theta)+\sqrt{3})=0 \]

OpenStudy (anonymous):

Ok? so? what do I do next?

OpenStudy (thomas5267):

What does it tell you about the value of \(\cos(\theta)\) and value of \(2\sin(\theta)+\sqrt{3}\). The product of the two has to be 0.

OpenStudy (anonymous):

Ok, I get that.

OpenStudy (anonymous):

But how do I find the actual values?

OpenStudy (anonymous):

Hello?